# Naturalistic Computational Cognitive Science Towards generalizable models and theories that capture the full range of natural behavior

### v2

@@ 4.4 Learning from naturalistic data allows us to ask new questions about the origins of knowledge
Example 1: language models and learnability. Recent progress in language modeling has suggested challenges to prior assumptions about language learnability and innateness. We briefly sketch these issues; see Piantadosi, (2023) and Futrell and Mahowald, (2025) for thorough accounts. Classical approaches to language processing focused on simple models and ideas (e.g. Chomsky,, 2014), but in doing so imposed strong assumptions. Under certain strong assumptions, Gold, (1967) proved that even relatively simple languages cannot be learned from input alone. This proof contributed to arguments that for innate universals such as recursion, and separation of syntax from semantics (Chomsky,, 1957, 1965, e.g.)—ideas that were subsequently influential in arguments about the nature of cognition more broadly (Fodor et al.,, 1975; Fodor and Pylyshyn,, 1988; Quilty-Dunn et al.,, 2023). However, these theories did not attempt to model language processing in all its naturalistic detail.
@@
By contrast, when deep-learning models learn from large amounts of naturalistic data, these models acquire aspects of both syntax (Manning et al.,, 2020) and semantics (Li et al., 2021a, )—and account for many behavioral and neuroscientific phenomena (Schrimpf et al.,, 2021). Although these models often learn from inhuman quantities of language (Wilcox et al.,, 2024), even models trained on a human-like quantity can learn complex syntactic features that were previously considered evidence for innate knowledge (e.g. Wilcox et al.,, 2023). Likewise, while some “impossible” languages that (adult) humans struggle to learn had been considered evidence for universal grammar, Kallini et al., (2024) show that language models can more easily learn real languages than impossible ones. Finally, controlled experiments—like those from Misra and Mahowald, (2024), reviewed above—begin to illustrate which features of experience contribute to syntactic generalization: particularly the importance of naturalistic variation. Thus, studying learnability from naturalistic data has helped to reshape our understanding of the necessary and sufficient features for acquiring linguistic capabilities.
@@
Example 2: naturalistic data can lead to brain-like functional specialization. Given the social importance of face perception, it might seem likely that face recognition is innately encoded in the brain. Indeed, the discovery of the Fusiform Face Area (FFA)—a visual region specialized for face perception (Kanwisher and Yovel,, 2006)—supports this hypothesis. However, recent computational work shows that this kind of specialization can emerge from a domain-general computer-vision model purely through training on a naturalistic distribution of object and face recognition tasks (Dobs et al.,, 2022). As the authors state: “It may be that the only inductive bias humans need to develop their face system is the already well-established early preference of infants to look at faces.” Of course, this does not necessarily imply that the FFA is not innately specified. However, it does illustrate how training computational models over naturalistic data can help us generate hypotheses about the causes of our neural organization.
5 Building generalizable models
Figure 6: Overview of key strategies cognitive science can adopt to develop generalizable models. (Figures reproduced from Geirhos et al.,, 2018; Bowers et al.,, 2023)
ML research focuses on developing generalizable models. While we focus on their utility for working with naturalistic tasks, we believe cognitive science can broadly benefit from improved practices for developing generalizable models. Here, we highlight three key practices of empirical ML research that we believe computational cognitive science can adopt. The first is frictionless reproducibility: developing research artifacts that can be reused and repurposed with minimal effort (§5.1). The second is a focus on developing generalizable models that can learn successfully across many datasets (§5.2). We argue that these two offer the foundation for the third: collectively collaborating on important research problems through dynamic benchmarks (§5.3).
Before we continue, we acknowledge that modern empirical ML is a young field which, like other fields, has faced challenges in rigor and reproducibility (Melis et al.,, 2017; Lucic et al.,, 2018; Riquelme et al.,, 2018; Henderson et al.,, 2018; Recht et al.,, 2019; Agarwal et al.,, 2021). Despite these challenges, we believe that some of the practices of ML are valuable for improving rigor and reproducibility, as we highlight below.
@@ 5.1 Frictionless reproducibility: a superpower
[…]
In recent years, data-centric sciences are experiencing a profound transition to what David Donoho calls “frictionless reproducibility” (Donoho,, 2024; cf. Recht,, 2024). Donoho identifies three key pillars to frictionless reproducibility: data sharing, code sharing, and competitive challenges. We focus on the first two in this section, and competitive challenges in §5.3. Critically, with frictionless reproducibility, the research artifacts enable future researchers to exactly re-execute the complete workflow with minimal effort. Thus, good ideas are spread, adopted, and improved with little friction. Donoho argues that ML is ostensibly the most successful at frictionless reproducibility. Below we review some strategies that may aid cognitive science.
@@
R1. Use standardized evaluation protocols where models are compared via a common interface. Machine learning’s foundation of standardized evaluation protocols traces back to 1959, when Highleyman and Kamentsky, (1959) created the first alphanumeric pattern-recognition dataset and the first ever “train-test” split. This strategy continued with the famous MNIST dataset (LeCun et al.,, 1998). MNIST simplified data-sharing and evaluation by pre-processing data, fixing train and test distributions, and sharing the data online. This standardization enabled fair comparisons between diverse approaches, from boosted stumps (Kégl and Busa-Fekete,, 2009) to support vector machines (Cristianini and Scholkopf,, 2002) or nearest neighbors (Khan,, 2017).
@@
Standardized evaluation protocols serve a critical function: mitigating personal bias in model assessment. The evolution to Caltech 101 (Fei-Fei et al.,, 2004) and ImageNet (Deng et al.,, 2009) demonstrates this principle, as increasingly naturalistic datasets enabled fair comparisons between diverse model families. This enabled AlexNet (Krizhevsky et al.,, 2012) to validate deep learning in an era when many researchers doubted its potential.
@@
Thus, we recommend that cognitive scientists establish clear hold-out data for both model development and data analysis. This is especially critical for confirming exploratory results (Poldrack et al.,, 2017). Combined with standardized evaluation protocols, this strategy enables fair comparison between diverse theoretical frameworks ranging from program synthesis to deep neural networks to reinforcement learning.
@@
R2. Make data accessible with a single line of code. Fully processed research data should be programmatically accessible with a single line of code, requiring no permissions. While uploading data to platforms like the Open Science Framework (OSF) is valuable, we additionally recommend platforms like Hugging Face that offer version control, free hosting, authentication-free access, access with one line of code, and uploading data with a few lines of code. This removes friction from reusing and updating datasets.
Model Citations Age GitHub Stars Single file examples
ResNet
He et al., (2016)
247,921 8 years 2,300 model
Contrastive Learning
He et al., (2020)
13,480 4 years 4,800 model, training
Object Discovery
Locatello et al., (2020)
790 4 years 394 model
RL Library
Huang et al., (2022)
252 2 years 5,600 model 1, model 2
RLHF Method
Rafailov et al., (2024)
1,730 1 year 2,200 train
Multi-agent RL
Rutherford et al., (2023)
24 0 years 433 model
Table 1: Comparison of ML libraries that adopted the single file philosophy.

### v4

@@ 4.3 Learning from naturalistic data allows us to ask new questions about the origins of knowledge
Example 1: language models and learnability. Recent language modeling works raise challenges to prior assumptions about language innateness. [^2] We briefly sketch these issues; see Piantadosi (2023) and Futrell and Mahowald (2025) for thorough accounts. Classical approaches to language focused on simple models (e.g. Chomsky, 2014) under strong assumptions. Using certain assumptions, Gold (1967) proved that even relatively simple languages cannot be learned from input alone. This contributed to arguments that language’s precursors must be innate (Chomsky, 1957, 1965, e.g.)—which influenced cognitive science writ large (Fodor and others, 1975; Fodor and Pylyshyn, 1988; Quilty-Dunn et al., 2023). However, these theories did not attempt to model language processing in all its naturalistic detail.
@@
By contrast, deep-learning models that learn from plentiful naturalistic data acquire aspects of syntax (Manning et al., 2020) and semantics (Li et al., 2021a)—and capture behavioral and neuroscientific phenomena (Schrimpf et al., 2021). Although these models often learn from inhuman quantities of language (Wilcox et al., 2024), even human-like quantities allow learning complex syntactic features that were previously considered evidence for innate structure (e.g. Wilcox et al., 2023). Furthermore, controlled experiments—like Misra and Mahowald (2024), reviewed above—show how naturalistic variation contributes to this syntactic generalization. Thus, studying learnability from naturalistic data has helped to reshape our understanding of the necessary and sufficient features for acquiring linguistic capabilities.
@@
Example 2: naturalistic data can lead to brain-like functional specialization. Deep neural networks trained on natural images have long predicted neural responses from V1 through IT (Olshausen and Field, 1996; Lee et al., 2007; Yamins and DiCarlo, 2016; Doshi and Konkle, 2023), establishing them as candidate models of visual processing. Recently, researchers have found that training with naturalistic data can yield brain-like functional specialization—something once thought to require innate hard-wiring. Given the social importance of faces, it might seem likely that face recognition is innately encoded in the brain. Indeed, the discovery of the Fusiform Face Area (FFA)—a visual region specialized for face perception (Kanwisher and Yovel, 2006)—supports this view. However, Dobs et al. (2022) showed that this specialization can emerge from a domain-general vision model trained on a naturalistic distribution of object and face recognition tasks. As the authors state: “It may be that the only inductive bias humans need to develop their face system is the already well-established early preference of infants to look at faces.” Of course, these results do not necessarily imply that the brain’s functional organization is not innately specified. However, they illustrate how training computational models on naturalistic data can generate new hypotheses about the origins of neural organization.
Example 3: naturalistic multi-view signals can yield human-level 3D shape perception. While large vision models trained on classification tasks capture some features of human-like vision, they show subhuman performance in areas like 3D object perception (Bonnen et al., 2024; cf. Malhotra et al., 2023). However, recently Bonnen et al. (2026) suggest that this may be due to these models lacking another key feature of human visual experience: the experience of seeing a scene from multiple angles (due to movement or stereo vision). Indeed, the authors show that models that are trained to predict view positions from multiple views of the same scene can match human-level 3D shape perception accuracy—thus showing how naturalistic learning signals beyond classification or image-only self-supervision might contribute to human 3D perceptual abilities.
5 Building generalizable models for naturalistic settings
through systematic and collective effort
How do we build generalizable models that predict and explain human behavior in naturalistic settings? ML has made substantial progress in building models that exhibit aspects of intelligence in settings closer and closer to the real world. We argue that this trajectory holds lessons for cognitive scientists who aim to model human-like intelligence in such settings.
The first lesson is that progress requires systematic iteration from simpler settings to more naturalistic ones. In principle, a single researcher can systematically cover the space of experiments needed to build generalizable models. In practice, the design space is vast. This makes individual progress difficult. The second lesson is that collective effort enables fast progress over a vast design space. Initially, ML researchers worked on disparate, self-defined problems, leading to diffuse progress (Recht, 2024). Progress accelerated when the field adopted shared benchmarks. A widely adopted benchmark is a sociological object: it certifies, often through peer review, that experts judge a problem worth solving. With the right infrastructure, this allows large populations of researchers to make fast progress together.
We propose that cognitive science adopt dynamic meta-benchmarks: living collections of benchmarks that researchers continually contribute to, where each benchmark tests a specific facet of a cognitive capacity (§5.1, see Figure 6). Consider object recognition. One facet might test whether classification relies on shape or texture; another whether shape processing uses local or global features; a third how classification interacts with scene context. Meta-benchmarks grow in two ways: researchers can increase the naturalism of existing facets (e.g., expanding shape-versus-texture stimuli from simple shapes to shapes derived from naturalistic stimuli; Geirhos et al., 2020), or add new facets exposing untested dimensions such as whether objects are classified by local- or global-shape features (Baker et al., 2020). In either way, benchmarks can cumulatively incorporate experimental tests of new hypotheses, without relinquishing the demand to explain prior results. A model that predicts human behavior across the full meta-benchmark provides evidence for a generalizable account of human capacities within that domain; failures on particular benchmarks identify where an account breaks down.
Adopting meta-benchmarks raises an engineering problem: how can we develop models that can operate across many benchmarks, stimulus sets, and task sets? The key is to develop models with an eye towards generalization by ensuring both that models generalize to new data, and that their development procedure generalizes as well (§5.2). That is, the procedure used to develop a model for one benchmark should allow development for other benchmarks with minimal modification. This is essential for efficiently validating models across a full meta-benchmark.
But how do we adopt development strategies that promote reuse across many benchmarks, stimulus sets, and task sets? One key strategy is frictionless reproducibility  5.3; Donoho, 2024), where researchers develop research artifacts that can be reused and repurposed with minimal effort. This includes standardized evaluation protocols and common interfaces that allow researchers to easily run any model against any benchmark. We argue that together, these pieces will enable cognitive scientists to develop generalizable models that capture the full range of natural behavior for a cognitive capacity in question.
Figure 6: How cognitive scientists can systematically build generalizable models of cognition that explain human behavior on naturalistic data. Our key proposal is leveraging dynamic meta-benchmarks5.1) where (a) individual benchmarks aggregate human data from experiments studying facets of a cognitive capacity. The benchmarks evolve as researchers add new facets—for example, extending object classification benchmarks to test whether classification relies on shape versus texture, then making stimuli (at least partially) reflect real-world data distributions. (b) Models are developed with generalization as the primary goal (§5.2): the same model should predict human behavior across all benchmark facets. (c) To facilitate frictionless reproducibility5.3), models are evaluated through a standardized interface. (d) This makes it easy for researchers to comprehensively evaluate their hypothesis (i.e. model) against a large swath of human behavioral data capturing different facets of a cognitive capacity. The best hypothesis is then the one that captures human behavior over the most human data. (Images reproduced from Geirhos et al., 2018)
5.1 Benchmarks: catalysts for progress and innovation
Since at least Newell (1994), cognitive science has sought to develop unified, task-performing models that explain cognition as a whole (Kriegeskorte and Douglas, 2018; Almaatouq et al., 2024). While cognitive scientists are aware that relying on a single paradigm may reduce the generalizability of conclusions (Holzmeister et al., 2024; Yarkoni, 2022), the field currently lacks infrastructure for this kind of multi-paradigm evaluation. The result is fragmentation and Newell’s critique—that piece-wise investigation of brain components would not yield a cohesive understanding (Newell, 1973)—remains relevant forty years later (Schrimpf et al., 2020; DiCarlo et al., 2023); findings from different experimental paradigms are difficult to build upon and resist integration, even within a domain (Almaatouq et al., 2024). We discuss how two concepts from ML—benchmarks and leaderboards—can enable collective, systematic effort within cognitive science.
Recommendations for cognitive science
R1. Embrace the leaderboard principle. A common concern with benchmarks is that researchers will “overfit” to test sets, which theoretically should underestimate the true test error (Hastie et al., 2009). Yet despite widespread iteration on benchmarks in machine learning the field has produced models with strong performance on held-out data. This apparent paradox is partially explained by a phenomenon known as the “leaderboard principle” (Hardt and Recht, 2022). Researchers typically only publish and build upon models that demonstrate substantial improvements over prior results, rather than minor variations. This selective pressure towards meaningful advances effectively constrains the degree of adaptation to test sets, as researchers focus on substantial improvements rather than exhaustively exploring the test set’s peculiarities. Nevertheless, a risk of overfitting to data remains. We argue that cognitive science can avoid this through the creation of dynamic phenomena-oriented benchmarks, where researchers can contribute new challenges that identify a missing component of a cognitive domain of interest. We detail this below.
R2. Develop dynamic benchmarks to iteratively capture important phenomena. ML benchmarks are not collected to test specific hypotheses (Hardt and Recht, 2022) and so may seem unscientific. However, cognitive science can embrace scientific benchmarks that allow testing specific hypotheses, while situating those tests within the context of the broader phenomena of the cognitive domain. We use the story of “Brain-Score” as an illustrative example of initial steps in this direction.
Brain-Score (Schrimpf et al., 2020) began as a benchmark collecting behavioral and neural data from humans and monkeys on object recognition tasks. One common critique of Brain-Score is that it does not test hypotheses about human object recognition. For example, Bowers et al. (2023) noted that matching human predictions is insufficient to identify a good model of human object recognition. The authors reference 9 experiments from the psychophysics literature which presented stimuli generated to test specific hypotheses—for example, testing use of local shape features (Baker et al., 2020), shape vs. texture  (Geirhos et al., 2018), and contours  (Puebla and Bowers, 2022)—and showed that deep neural networks failed to capture human behavior across these. In response, DiCarlo et al. (2023) added many of these stimulus sets to Brain-Score, so that future models can be evaluated against all of them. This story illustrates what an exciting strand of computational cognitive science research could look like: one where researchers develop models and experiments that easily integrate with prior findings on a phenomenon thanks to phenomena-oriented benchmarks, generalizable models (§5.2), and frictionless reproducibility (§5.3).
We are not the first to identify that cognitive science is fragmented and to suggest benchmarks that aggregate results from the literature as a solution. Indeed, researchers spanning memory, timing, confidence and more have already begun to do so (MacWhinney and Snow, 1985; Oberauer et al., 2018; Aydoğan et al., 2024; Rahnev et al., 2020; Frank et al., 2017). Here, we both encourage more subdisciplines within cognitive science to follow suit, and emphasize the importance of frictionless reproducibility and building generalizable models. While prior work has begun developing benchmarks, documentation for downloading and using the data is sometimes lacking; download access is not always authentication-free; and maybe most importantly, if one wants to add new data, there is not always an easy way to re-run all previous models and ensure that findings are commensurate. More critically, while prior benchmarks have focused on aggregating experimental results from “simplified” experiments, we advocate for dynamic benchmarks whose development is oriented around explaining data from increasingly naturalistic experiments (i.e. experiments with either parametrically generated naturalistic stimuli or, eventually, stimuli taken from real-world conditions).
5.2 Developing models with an eye towards generalizability
Ideally, the models we develop generalize beyond the setting they were designed for. This requires two things. First, that the model is generalizable, i.e. can learn from, operate over, and perform new tasks. Second, that the behavioral (and potentially neural) predictions made by our models capture data on these new tasks. Clearly, the first requirement is a prerequisite for the second. Here, we detail suggestions from ML for how cognitive science can develop more generalizable models.
Recommendations for cognitive science
R3. Iterate between (1) model-design (2) task design where the other is fixed. One crucial lesson from ML research is the importance of separating model development from task design. Patterson et al. (2023) strongly caution against simultaneously developing both the problem setting (datasets or tasks) and the solution method, as researchers may inadvertently design evaluations that favor their proposed solution. Historical examples illustrate this risk. When developing Q-learning algorithms, researchers modified the classic pendulum control problem by adding episode cutoffs and random state resets (Machado et al., 2018). These modifications made Q-learning more tractable by implicitly improving exploration, but thereby obscured fundamental limitations in Q-learning’s exploration compared to policy gradient methods. Maintaining a strict separation between task design and model development enforces a clear delineation between hypothesis formation and testing, which is important for reducing experimenter bias. This is especially important because researchers tend to find “positive” results (Scheel et al., 2021; Sarafoglou et al., 2022), a bias exacerbated by an increasingly competitive academic landscape (Lee, 2014; Reithmeier et al., 2019; Woolston, 2021). Much of ML’s progress can be attributed to researchers evaluating their models on independently developed datasets and tasks—a practice that both simplifies the research process and promotes scientific rigor.
Thus, we recommend that cognitive science also develop models on tasks/stimuli developed by others. This may seem to contradict the scientific method, where researchers craft experiments to test their specific hypotheses. The data from prior work was for a different hypothesis! However, this is precisely where dynamic, phenomena-oriented benchmarks can offer a solution. First, where researchers hypothesize a failure of existing theories or models, they can develop a new experimental paradigm that illustrates it, and then commit this paradigm to a dynamic benchmark. Once the paradigm has been publicly committed, they can move on to proposing and testing new models to explain it—which will simultaneously be tested for their continued ability to explain the other phenomena included in the benchmark. In this way, researchers can pursue hypothesis-driven experiment- and model-development, while maintaining scientific rigor by avoiding the temptation to overfit tasks and models to each other. We detail the practical approaches that enable this beneficial cycle further in the next section.
R4. Develop models with simplified tasks, but ensure they work with increasingly naturalistic tasks. The frame problem (Shanahan, 2004) and the “simulation is doomed to succeed” critique (Grim et al., 2013) flag a recurring failure: the evaluation tasks we design rarely capture the full complexity of the phenomenon we care about. ML’s response is to demand that models succeed in both simplified and increasingly complex settings. RL exemplifies the trajectory: initial environments were grid-worlds (Sutton and Barto, 2018), later 2D video games (Bellemare et al., 2013), then virtual 3D homes (Szot et al., 2021; Li et al., 2021b), and now open-world environments (Matthews et al., 2024; Hughes et al., 2024). Even with this push toward naturalism, researchers still rely on simpler settings to study specific aspects of model behavior and generalization (Osband et al., 2019; Obando Ceron et al., 2024). Indeed, the standard strategy is to first develop a model in simplified settings, then verify that it continues to work in more naturalistic ones. This has yielded both theoretical (Grimm et al., 2020; Lyle et al., 2022) and empirical (Obando Ceron et al., 2024; Farebrother et al., 2024) understanding of how RL models work, alongside performant models for complex settings—beating world experts in Go (Silver et al., 2017) and controlling real robots (Levine et al., 2016; Cheng et al., 2024).
R5. Develop models that work on many tasks. The ultimate goal of ML is generalization (Highleyman and Kamentsky, 1959; Hardt and Recht, 2022; Recht, 2024). Initially, this meant models that generalize to new data. Now, it means both models and learning procedures that generalize to new settings. Thus, evaluating models across diverse datasets has become foundational in ML. Early examples include “dropout” (Srivastava et al., 2014), which demonstrated improvements across 6 image and speech datasets, and MoCo (He et al., 2020), which validated its approach on 9 vision datasets. Another notable example is the seminal “Deep Q-Network” paper (Mnih et al., 2015), which showed that a single neural network architecture with fixed hyperparameters [^3] E.g., the number of layers or the learning rate. could master 50 different Atari games. The field has since evolved to organize environments by the cognitive capabilities they test (e.g., exploration, generalization, manipulation) (Patterson et al., 2023), with researchers developing models on some task sets and validating on entirely different ones to ensure robust generalization (Machado et al., 2018).
We recommend cognitive scientists also test the models that support their theories more broadly. Importantly, we should evaluate models on different tasks from those we use to develop them (Machado et al., 2018). This further helps prevent our model (and thereby theory) from “overfitting” to a particular paradigm.
@@ 5.3 Frictionless reproducibility: a superpower
Cognitive science and psychology face substantial friction reusing prior data and code (Nosek et al., 2015; Hardwicke et al., 2018): data are rarely standardized, analyses are often underreported, and shared code frequently contains errors (Poldrack et al., 2017). In a large-scale replication analysis, $38\%$ of code was not usable and only $31\%$ of workflows reproduced (Hardwicke et al., 2018). Meanwhile, data-centric sciences are undergoing a profound transition—what Donoho (2024) calls frictionless reproducibility—in which good data- and code-sharing practices combine with competitive challenges to let researchers re-execute prior workflows with minimal effort, accelerating progress dramatically. ML, Donoho argues, is ostensibly the most successful at frictionless reproducibility. Below we review some strategies that cognitive science can adopt. §5.1 addresses competitive challenges via dynamic meta-benchmarks; here we focus on data and code sharing.
@@
R6. Use standardized evaluation protocols where models are compared via a common interface. Machine learning’s foundation of standardized evaluation protocols traces back to 1959, when Highleyman and Kamentsky (1959) created the first alphanumeric pattern-recognition dataset and the first ever “train-test” split. This strategy continued with the famous MNIST dataset (LeCun et al., 1998). MNIST simplified data-sharing and evaluation by pre-processing data, fixing train and test distributions, and sharing the data online. This standardization enabled fair comparisons between diverse approaches, from boosted stumps (Kégl and Busa-Fekete, 2009) to support vector machines (Cristianini and Scholkopf, 2002) or nearest neighbors (Khan, 2017).
@@
Standardized evaluation protocols serve a critical function: mitigating personal bias in model assessment. The evolution to Caltech 101 (Fei-Fei et al., 2004) and ImageNet (Deng et al., 2009) demonstrates this principle, as increasingly naturalistic datasets enabled fair comparisons between diverse model families. This enabled AlexNet (Krizhevsky et al., 2012) to validate deep learning in an era when many researchers doubted its potential.
@@
Thus, we recommend that cognitive scientists establish clear hold-out data for both model development and data analysis. This is especially critical for confirming exploratory results (Poldrack et al., 2017). Combined with standardized evaluation protocols, this strategy enables fair comparison between diverse theoretical frameworks ranging from program synthesis to deep neural networks to reinforcement learning.
@@
R7. Make data accessible with a single line of code. Fully processed research data should be programmatically accessible with a single line of code, requiring no permissions. While uploading data to platforms like the Open Science Framework (OSF) is valuable, we additionally recommend platforms like Hugging Face that offer version control, free hosting, authentication-free access, access with one line of code, and uploading data with a few lines of code. This removes friction from reusing and updating datasets.